A medical researcher is studying the effects of a drug on blood pressure. Subjects in the study have their blood pressure taken at the beginning of the study. After being on the medication for 4 weeks, their blood pressure is taken again. The change in blood pressure is recorded and used in doing the hypothesis test.

Change: Final Blood Pressure - Initial Blood Pressure

The researcher wants to know if there is evidence that the drug increases blood pressure. At the end of 4 weeks, 33 subjects in the study had an average change in blood pressure of 2.7 with a standard deviation of 5.4.

Find the p-value for the hypothesis test.

Your answer should be rounded to 4 decimal places.
Answer #1

Solution :

Given that,

Population mean = \mu = 4

Sample mean = \bar x = 2.7

Sample standard deviation = s = 5.4

Sample size = n = 33

This a right (One) tailed test.

The null and alternative hypothesis is,  

Ho: \mu \leq 4

Ha: \mu > 4

The test statistics,

t = ( \bar x - \mu )/ (s/\sqrt{n})

= ( 2.7 - 4 ) / ( 5.4 /\sqrt{}33 )

= -1.383

df = n - 1 = 32

P- Value = 0.9119